OpenStudy (anonymous):

Find the Sum (2a^2 - 3a +1) + (3a^2 + 2a -2) = x

OpenStudy (saifoo.khan):

5a^2-a-1 = x

OpenStudy (anonymous):

THANK YOU!

OpenStudy (saifoo.khan):

uw!

OpenStudy (anonymous):

(2a^2 - 3a +1) + (3a^2 + 2a -2) = x 5a^2-a-1=x, we can solve for a(quadratic equation) when x=0